the sum of three numbers is 95. the sum of the first and the second is twenty five more than the third number and the third is twice the first. find the numbers.
Answer by ankor@dixie-net.com(12684)

Three numbers: x, y, z
;
"the sum of three numbers is 95."
x + y + z = 95
;
" the sum of the first and the second is twenty five more than the third number "
x + y = z + 25
:
" the third is twice the first."
z = 2x
:
find the numbers.
:
rewrite the 2nd equation and subtract from the 1st equaion
x + y + z = 95
x + y - z = 25
-----------------subtraction eliminates x and y, find z
2z = 70
z = 35
therefore
2x = 35
x = 17.5
:
Find y:
17.5 + y = 35 + 25
y = 60 - 17.5
y = 42.5
;
:
Check the sum of the three numbers
17.5 + 42.5 + 35 = 95